A 1.4-kg object is pulled horizontally along the floor against 3.2 N of kinetic friction. If the object accelerates at 5.8 m/s2, what is the value of the applied force?
1
Expert's answer
2019-10-24T09:28:41-0400
"ma=F-F_f"

"F=ma+F_f=1.4(5.8)+3.2=11.32\ N"